OpenStudy (deoxna):

Is there anyway to relate momentum or kinetic energy to force? I remember a while ago I had a physics problem about a rifle shooting out a bullet at a given velocity, and then we had to calculate the momentum of the rifle's recoil. However, is the "damage" caused to the person holding the rifle because of this momentum, or is there a way of calculating the total force exerted on the person?

OpenStudy (anonymous):

force=change in momentum/time

OpenStudy (anonymous):

That's correct, yeah\[\vec F = \frac{d \vec p}{dt}\]Also note that \[E_{kin} = \frac{m \, \vec v^2}{2} = \frac{\vec p^2}{2m}\]